### v2

# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

Abstract.
Humans think visually—we remember in images, dream in pictures, and use visual metaphors to communicate. Yet, most creative writing tools remain text-centric, limiting how authors planing and translating ideas. We present Vistoria, a system for synchronized text–image co-editing in fictional story writing that treats visuals and text as co-equal narrative materials. A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and annotations serve as essential instruments for ideation and organization. Drawing on theories of Instrumental Interaction and Structural Mapping, Vistoria introduces multimodal operations-lasso, collage, filters, and perspective shifts that enable seamless narrative exploration across modalities. A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, enabling writers to explore divergent directions, embrace serendipitous randomness, and trace evolving storylines. While multimodality increased cognitive demand, participants reported stronger senses of authorship and agency. These findings demonstrate how multimodal co-editing expands creative potential by balancing abstraction and concreteness in narrative development.
Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interation, Structure Mapping, Direct Manipulation
journalyear: 2026copyright: acmlicensedconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Creant el demà junts, Barcelonabooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Creant el demà junts, Barcelona
Figure 1. Vistoria supports a cyclic workflow in which multimodal artifacts and ideas co-evolve. (A) Instrument Interaction: a unified set of tools—Collage, Lasso, Filter, and Perspective Shift—enables coupled editing of image and text to produce image–text “cards.” (B) Artifacts → Idea: the resulting visual and textual artifacts stimulate new concepts and story directions. (C) Cognitive Support: a structural mapping view aligns image and text variations, making alternatives comparable and keeping narrative structure consistent. (D) Ideas → Artifacts: emerging ideas are materialized back into new cards, closing the loop and driving iterative ideation, exploration, and integration.
1. Introduction
Human cognition is deeply rooted in visual processing (Rezaee, 2011; Huang et al., 2024; Kanellopoulou et al., 2019). We recall experiences as spatial scenes, build mental models through imagery, and rely on techniques such as memory palaces to organize abstract information. Everyday reasoning often involves manipulating internal visual representations—planning a route, rearranging objects in mind, or picturing alternative outcomes. Language use is similarly tied to imagery—comprehension frequently evokes mental pictures, and abstract ideas are often conveyed through spatial metaphors such as path, framework, or perspective (Meng et al., 2023). This aligns with Dual Coding Theory, which argues that cognition integrates both verbal and nonverbal channels, with distinct representational capacities, to support imagination, reasoning, and communication (Clark and Paivio, 1991).
Given this inherent multimodality of thought, a critical question arises: why do our most sophisticated creative practices, such as narrative writing, remain confined to unimodal tools? Fictional story writers often describe richly visualized scenes and manipulate complex spatial relationships in their imagination, yet must translate these processes through the narrow channel of sequential text (Mukramah et al., 2023; Listyani, 2019; Rao et al., 2024; Wang et al., 2025c). This mismatch between the multimodal nature of creative cognition and the unimodal design of creative tools constrains expressive potential (Zhao et al., 2025; Lee et al., 2024).
In fictional story writing, the most demanding phases are planning and translating (Gero et al., 2022; Flower and Hayes, 1981). Authors must organize imagined scenes, characters, and events into coherent structures, then render these abstractions into linear prose (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023). Visuals can scaffold these processes by stimulating ideas, supporting organization, and enriching narrative detail (Mukramah et al., 2023; Listyani, 2019; Ali and Parikh, 2021). However, traditional visual aids have been limited in scope and scalability, offering only passive inspiration rather than dynamic, interactive support for the iterative writing process (Kreminski et al., 2024; Choi et al., 2024; Yan et al., 2023; Mukramah et al., 2023; Listyani, 2019). Large language models (LLMs) have unlocked unprecedented generative capabilities, enabling rapid creation of diverse narrative elements through combined text and visual imagery at previously unattainable scales (Fu et al., 2024). LLM-powered systems have explored visual integration through character construction tools (Qin et al., 2024), visual analytics for branching storylines (Mishra et al., 2025; Wang et al., 2025a), script visualization (Rao et al., 2024), diagram-prose synchronization (Masson et al., 2024b), and text-editing metaphors from image manipulation (Masson et al., 2024a).
However, the use of visuals remains peripheral across these approaches. They are often treated as prompts, overlays, or loosely coupled representations with unidirectional interaction. Our formative Wizard-of-Oz (WOz) co-design study with 10 writers revealed that sketches, annotations, and image manipulations are essential instruments for externalizing mental images and organizing narrative ideas. What is missing from existing LLM-powered creativity tools for fictional story writing is a unified multimodal substrate that directly supports iterative oscillation between planning and translating, where text and visuals serve as co-equal, manipulable materials in a bidirectional co-editing loop.
To bridge this gap, we developed Vistoria, a system that transforms fictional story writing from a unimodal text process into a multimodal co-editing experience where synchronized text-image manipulations through polymorphic operations—lasso, collage, perspective shifts, and filters. The design enables writers to fluidly navigate between abstract conceptualization and concrete visualization while creating a unified multimodal substrate for narrative development. These operations enable writers to reorganize narrative elements, explore alternative viewpoints, and refine story details through a bidirectional loop where text and visuals continuously inform one another. The design of Vistoria was grounded in two key theoretical frameworks. Instrumental Interaction guides consistent operations across modalities to reduce switching costs and sustain flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025), while Structural Mapping (Gentner, 1983; Norman, 2013) explains how verbal and visual representations can be aligned so edits in one modality inform the other.
We conducted a controlled study with 12 participants to evaluate Vistoria’s effectiveness in supporting multimodal narrative creation. The results indicate that participants successfully integrated Vistoria’s synchronized text-image co-editing into their creative workflows, employing four distinct multimodal strategies to enhance storytelling. Participants demonstrated enhanced creative expression by using images as catalysts for richer descriptions, leveraging visual editing operations to drive narrative pivots, and employing multimodal elements to externalize and interconnect story branches. The cross-modal dialogue facilitated creativity through divergent exploration of multiple narrative directions, serendipitous discoveries from AI-generated content, and culture-driven associations that activated personal memories and references. While multimodality increased cognitive load, study participants reported heightened authorship, agency, and immersion, characterizing the system as a collaborative partner that expanded their imagination. Ultimately, creativity emerged from the dynamic interplay between modalities, where text and imagery continuously shaped one another.
In summary, this work contributes:
  • A WOz co-design study with 10 writers examined the practices and needs of using multimodal elements (text, sketches, and visual imagery) to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
  • Vistoria, a multimodal co-editing system that unifies text and visual images through polymorphic operations and a bidirectional editing loop to support planning and translating in fictional story writing;
  • A controlled usability study with 12 participants demonstrates the effectiveness of Vistoria, showing that multimodal co-editing enhances expressiveness, coherence, and authorial agency in fictional story writing.
2. Related Work
2.1. Using Visuals to Support the Cognitive Process of Fictional Story Writing
Fictional story writing is distinct from argumentative or expository genres in its emphasis on imagination, world-building, and character development (Doyle, 1998). Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, non-verbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
The Cognitive Process Model of Writing (Gero et al., 2022; Flower and Hayes, 1981) frames writing as recursive processes of planning, translating, and reviewing. In fictional story writing, the transition from planning to translating is especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load. However, visual representations can scaffold this process by externalizing abstract ideas. Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019). In practice, sketches, maps, and diagrams help structure plot, setting, and character relationships during planning, and serve as cognitive anchors during translating to support coherence and consistency (Zhao et al., 2025). Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and imagery systems function separately but also interact, creating richer memory traces when information is encoded in both modalities. In fictional writing, visual representations of narrative elements complement verbal planning, making abstract concepts more concrete and retrievable. When writers encounter difficulties in translation, visual anchors provide alternative access to imaginative content, reducing cognitive load and enabling more fluid expression (Barsalou, 2008).
Building on this foundation, a variety of creativity support tools have incorporated visuals into different phases of writing. Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, often through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Long et al., 2024; Qin et al., 2024). Translation-focused tools like ScriptViz and Script2Screen aim to align textual composition with visual referents, either by retrieving reference visuals from movie databases (Rao et al., 2024) or by synchronizing scriptwriting with audiovisual scene creation (Wang et al., 2025c). Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025a), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
Despite these advances, visuals in prior systems remain peripheral: they serve primarily as references, scaffolds, or analytic overlays, rather than as co-equal, manipulable materials. This treatment constrains multimodal creative expression by limiting writers to either text-first workflows or loosely coupled visual aids (Zhao et al., 2025). To bridge this gap, our work introduces a system that treats visuals and text as fully integrated, co-editable artifacts, directly supporting the planning and translating phases of fictional story writing so that imagination can be externalized and refined seamlessly across both modalities.
2.2. LLM-powered Multimodality Tools for Creativity in Content Creation
Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2024a). Powered by advanced LLMs, these systems address fundamental barriers through three complementary mechanisms. First, externalizing creative structures. Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) turn abstract intentions into manipulable interface objects or reusable design tokens, making spatial, stylistic, and compositional relations tangible. Second, reducing cognitive load with multimodal input: DrawTalking (Rosenberg et al., 2024) and Code Shaping (Yen et al., 2025) distribute expression across sketch, voice, or annotations, lowering the reliance on precise verbal specification. Third, supporting iterative refinement. Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025b) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references. However, most current systems remain fragmented: they either support isolated stages (e.g., ideation, editing) or couple modalities only loosely, often within a one-shot prompting paradigm that underutilizes LLMs’ potential for sustained dialogue and refinement (Dang et al., 2023).
Seeking tighter coupling for fictional story writing, recent systems push integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024). WorldSmith grounds worldbuilding in text–sketch edits and region-based filling to support users iteratively visualize and modify elements of their fictional world (Dang et al., 2023). XCreation (Yan et al., 2023) supports cross-modal storybook creation by integrating an interpretable entity-relation graph, improving the usability of the underlying generative structures. Toyteller (Chung et al., 2025) maps symbolic motions to character actions. Visual Writing allows writers co-edit stories by directly manipulating visual representations of entities, actions, and events alongside text, enabling easier navigation, editing, and exploration of narrative variations (Masson et al., 2024b). These systems highlight the transformative potential of LLMs to couple narrative and visual modalities, yet they still lack a unified, synchronized substrate that supports fictional story writing as a continuous, multimodal workflow. Our work advances this direction by introducing such a substrate, where LLM-generated texts and visuals act as co-editable inspirational materials for iterative, coherent story development.
2.3. Instrumental Interaction and Structural Mapping Theory
Instrumental Interaction is central to understanding how users control and refine digital systems. Beaudouin-Lafon (Beaudouin-Lafon, 2000) proposed it as a shift from designing static interface elements to designing instruments that mediate between users and domain objects. A key principle is reification, which transforms abstract commands into persistent, manipulable objects (Kent, 2000). In computing, this elevates implicit system descriptions into explicit first-class entities. In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2023) and in Textoshop’s reification of abstract text editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2024a). A second principle is polymorphism, where the same instrument applies across contexts (Masson, 2023). This reduces cognitive load by enabling predictable, transferable patterns, e.g., copy–paste works consistently across text, images, and files (Angelini et al., 2015), and scrollbars operate similarly across documents, spreadsheets, and browsers (Masson, 2023). Finally, reuse allows users to replay or adapt prior operations, from macros to redo commands (Riche et al., 2025). Systems like Spacetime exemplify this by objectifying space, time, and actions into persistent containers, enabling edits to be carried forward as manipulable entities (Xia et al., 2018). Together, these principles reduce cognitive burden by externalizing interaction histories, making them manipulable, transferable, and extensible.
In our system, we extend this perspective to AI support for fictional story writing, where imagination must be continuously externalized and revised. We design interaction instruments that reify narrative elements into manipulable objects and, through polymorphism, let the same tools operate across LLM-generated text and images. This integration of reification and polymorphism enables writers to fluidly adjust multimodal outputs as persistent, tangible entities.
Structural Mapping Theory (SMT) is central to understanding how users interpret and compare complex representations. Gentner (Gentner, 1983) proposed SMT to describe how people form abstractions through analogical alignment of relational structures. Rather than relying on surface similarities, SMT highlights the role of alignable differences, which reveal systematic variations and help learners construct new knowledge. Building on this foundation, SMT-informed design emphasizes preserving object structure rather than abstracting it away. At the collection level, showing complete objects allows users to perceive relationships across the dataset (Yan et al., 2022; Wu et al., 2020). At the part level, systems can highlight structural correspondences by surfacing identity matches (literal overlaps) (Glassman et al., 2015), property matches (shared attributes) (Jasim et al., 2022), and conceptual matches (semantic commonalities) (Glassman et al., 2018). These correspondences support the iterative process of aligning, re-aligning, and refining mental models. Prior visualization and learning systems demonstrate the benefits of making such correspondences explicit, between code solutions (Glassman et al., 2015), textual variants (Glassman et al., 2018), or large language model outputs (Arawjo et al., 2024; Gero et al., 2024), helping users notice critical differences, generalize across examples, and avoid disengagement when facing “walls of text.” SMT thus serves as both a cognitive theory and a design constant for interactive systems that deepen users’ reasoning about collections of related objects.
In our system, SMT informs the design of synchronized text–image co-editing: textual fragments are explicitly mapped to visual elements based on their shared properties, matched concepts, or references to the same object, so edits in one modality lead to structurally aligned updates in the other, allowing users to compare, reorganize, and refine story elements across modalities.
3. Formative Study
Previous research shows that multimodal tools enhance fictional story writing by making abstract concepts tangible, reducing cognitive load, and improving creativity and coherence (Chung et al., 2022; Chung and Kreminski, 2024; Dang et al., 2023; Zhao et al., 2025). However, current tools treat images and texts as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types.
To address this gap, we conducted a Wizard-of-Oz co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981). Our investigation focused on three questions: (1) Multimodal information use: what types of multimodal content users employ and how they leverage these materials for idea generation; (2) Iteration and integration: how creators refine and combine multimodal artifacts in world-building and narrative development; and (3) Organization of inspirations: how creators organize, connect, and refine dispersed inspirations through multimodal manipulation. The WoZ setup simulated AI-assisted visual and textual support while sustaining the impression of an intelligent, interactive system.
3.1. Process
We designed a Wizard-of-Oz (WoZ) co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
3.1.1. Participants
We recruited 10 participants, each with an two or more years of creative writing experience. The group included three fictional story writers, three animation scriptwriters, two visual film creators, one new media creator, and one online fiction writer. Eight participants held a master’s degree or higher, and two held a bachelor’s degree.
Experimental Setup Three days prior to the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development. The 90-minute main session took place in Figma(Figma, 2025), which provided a collaborative canvas for free sketching, arranging text–image materials, and treating multimodal inputs as manipulable artifacts.
Session Process During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with a simulated AI assistant operated by the researcher (wizard). Rather than working toward a fixed output, participants iteratively developed stories of about 300 words while envisioning how the tool itself should behave. They freely requested content through voice, text, or hand-drawn sketches, and called upon the assistant for image and text editing as new needs emerged. The wizard responded in situ to simulate intelligent system behaviors, encouraging naturalistic exploration. To support this process, the experimenter used Claude (Anthropic, 2025) for text generation (drawing on outlines and randomness to stimulate ideation) and Midjourney (Midjourney, 2025) and GPT-4o (OpenAI, 2024) for image generation from prompts, sketches, and editing instructions. Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
3.2. Formative Study Findings
3.2.1. Using multimodal input to reify vague ideas
As shown in the Table LABEL:tab:formative_grouped, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
Sketches externalized vague intentions and spatial imagination. For example, P4 envisioned a scene where clown JoJo appeared on stage and created a sketch showing “JoJo emerging from a giant spinning wheel” with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details. Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories while clearly communicating their intentions. Participants frequently used textual annotations to specify story parts or image types for AI generation and guide generation direction. Images communicated style and mood expectations. P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “if possible, I want to use a ’supporting image’–a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.”
Table 1. Results of user strategies for manipulating multimodal elements in the WOz co-design study.
Stage Observed behavior User need / Insight N Interaction Example
Multi-modal input Text Generated prose should match the established world while allowing the injection of new elements 83 [Uncaptioned image] [Uncaptioned image]
Multi-modal input Sketches Make spatial relations/composition concrete 8 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated images Precisely locate inspiration from images 90 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated text Filter usable bits from many generations and reuse them 90 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Text notes / annotations Externalize temporary ideas for re-generation or final writing 30 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Collage elements Recompose fragments from cross-image to form new scenes 12 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Link elements Structure relationships between character / scene / object to connect the plot 40 [Uncaptioned image] [Uncaptioned image]
Translating Reconfigure elements in canvas Reorder scattered ideas by role / time / space into place for global understanding 4 [Uncaptioned image] [Uncaptioned image]
Translating Text (integration) Consolidate AI-generated text and image-inspired content into the draft 10 [Uncaptioned image] [Uncaptioned image]
3.2.2. Text–Image Interplay as Complementary Design Moves
Participants most often sought AI elaboration on characters, objects, and scenes, expressing the need for assistance with character design refinement, setting depictions, or object visualization. Combining image outputs with textual descriptions helped participants enrich their limited knowledge. For example, P7 requested AI-generated designs of an ancient Chinese poison bottle as a narrative element, noting her vague understanding of the concept.
These findings highlight the value of systems that accept heterogeneous multimodal input and help co-designers transform nascent ideas into concrete narrative materials.
Participants perceived text and images as distinct yet complementary in their creative writing. We observed an oscillation between abstraction (text) and concreteness (images) as a recurring co-design pattern, which reveals design opportunities for multimodal authoring tools.
Text as open imagination. Participants described text as a “blank canvas” for associative and structural thinking. P4 noted, “text allows me to imagine many things in my mind,” and P1 emphasized its “infinite imagination on a blank page.” P8 highlighted that text helped set up narrative structure before layering in visuals. This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than prematurely resolved.
Images as concreteness, inspiration, and feedback. Images grounded abstract ideas, while their randomness often sparked unexpected inspiration. P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P4 refined Lily’s behavior based on an unexpected visual detail, and P3 used images as feedback for progressive refinement. P10 likened it to a “look-and-write exercise” that scaffolds scene construction. These accounts highlight the value of image outputs not just as illustrations but as provocations. Designers can leverage it through image-based iteration, selection, and reinterpretation.
Complementary interplay. Participants emphasized that neither modality sufficed alone: images “set the vibe,” while text reframed meaning. P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence. This interplay suggests design opportunities for systems that orchestrate smooth transitions between modalities, enabling users to fluidly oscillate between abstract exploration and concrete elaboration.
3.2.3. Direct Manipulation of Multimodal Artifacts
Participants expressed a strong interest in treating text and images as manipulable, recombinable design materials. Two recurring practices pointed to design needs for more fluid multimodal manipulation.
Collaging and Recombination. Participants frequently merged elements across outputs to spark new ideas. P9 envisioned combining “the house from the first AI-generated images with the street from the second picture” to construct scenes, while P7 highlighted that “randomly combining characters and scenes” could inspire unexpected connections when accompanied by textual descriptions. As she explained, “if I can see an AI-generated image of my protagonist in one of the scenes, it helps me better imagine potential connections between elements that might otherwise seem unrelated.” Such practices illustrate the potential of collage and recombination as creative strategies.
Granular Editing and Annotation. Beyond recombination, participants desired fine-grained control over outputs. Sticky notes captured details for iteration and served as prompts for later development. Participants also wanted more localized operations, such as regenerating specific regions (P1, P3), extracting and reusing circled image elements (P5), or annotating character personas for refinement (P2). They left narrative prompts for later translation, e.g., P2’s note “Ending could be related to why this postman job even exists.” These behaviors emphasize annotation as both a vehicle for iteration and a bridge to subsequent writing.
Together, these findings suggest systems should enable flexible recombination, localized editing, and traceable annotations to help creators turn multimodal inspiration into actionable writing material.
3.2.4. From Fragmented Inspirations to Coherent Storylines
While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge. As P2 noted, “everything quickly became too messy on the canvas”, and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts. Furthermore, P6 wished for mind map–like tools to scaffold this process. Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” for ensuring cross-chapter consistency and avoiding logic conflicts (P4, P7).
These challenges point to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
3.3. Design Goals
Drawing on insights from our WoZ co-design study, prior work on multimodal LLM tools, and theories of Structural Mapping and Instrumental Interaction (Section 2.3), we identify four design goals for a multimodal content creation interface that supports the planning and translating phase in fictional story writing (Gero et al., 2022; Flower and Hayes, 1981; Zhao et al., 2025).
  • DG1: Reifying Multimodal Intentions into Manipulable Artifacts. Informed by findings in Section 3.2.1, our system should reify multimodal inputs (text, sketches, images) into first-class artifacts that can be annotated, rearranged, and recombined. Drawing on Instrumental Interaction, such reification should preserve fleeting intentions as reusable materials for later iteration; following SMT, these artifacts should function as alignable units that support comparison, linkage, and cross-modal reorganization.
  • DG2: Aligning Text and Images for Iterative Creative Exploration. Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, image refinements can inform descriptions, and alternative versions should make differences explicit. Grounded in Dual Coding Theory and SMT, this alignment should support cycles of divergence and convergence by allowing users to oscillate between abstraction (text) and concreteness (images).
  • DG3: Enabling Polymorphic Cross-Modal Manipulation. Informed by findings in Section 3.2.3, our system should support direct manipulation interactions that apply uniformly across modalities, guided by Instrumental Interaction’s principle of polymorphism. This uniformity should reduce switching costs and empower writers to fluidly manipulate and edit textual and visual fragments while maintaining narrative coherence.
  • DG4: Organizing and Reusing Fragments into Coherent Narratives. Informed by findings in Section 3.2.4, our system should cluster and organize fragments and fleeting ideations, surface hidden connections, and consolidate dispersed inspirations into coherent, evolving narratives. It should further support the reuse of narrative units (e.g., characters, settings, objects) to ensure consistency and foster pathways for story progression.
4. Vistoria System
In this section, we present the key features of Vistoria. As shown in Figure 2, the interface comprises three primary components: a left text editor, a central collapsible cluster panel, and a right canvas interface. The text editor houses the current story draft, serving as contextual information for content generation. The right canvas supports freeform sketching, text input, and comprehensive editing tools. The central cluster panel aggregates highlights and annotations from canvas content, organizing them by characters, objects, and scenes for reference and summary.
Together, these components create a workflow where narrative text, visual artifacts, and organizational structures remain continuously aligned, enabling writers to externalize, manipulate, and recombine ideas and materials across modalities.
4.1. Key Features
Figure 2. Vistoria’s interface: a left text editor, a central Cluster panel(can be collapsed when not used), and a right free-form Canvas. Cards (image-text pairs with main object keywords) are generated from user intentions and serve as the basic units; two modes—Exact Craft and Creative Spark—support precise realization vs. exploratory variation.
4.1.1. Reifying Intention through Multimodal Generation
Writers often externalize early intentions through sketches, short notes, or reference images, but these expressions are typically fleeting and difficult to refine into coherent narrative materials. Supporting the preservation and refinement of such multimodal intentions is therefore essential for sustaining ideation (DG1). To address this, Vistoria reifies multimodal inputs into persistent Cards that bundle an image, a story fragment, and extracted narrative objects (characters, objects, scenes). Drawing on the principle of Reuse, Vistoria reifies sketches and text from ephemeral marks into manipulable and reusable artifacts that remain on the canvas and can be invoked again as elements for regeneration, enabling writers to carry forward and iteratively develop their creative intentions. At the same time, we treat cards as alignable units: text and their corresponding images are designed to carry the same underlying meaning, with images serving as visualizations of the narrative described in text.
Vistoria further balances precision and exploration by offering two complementary generation modes (Figure 3). In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas. In Creative Spark mode, three cards were generate represent diverse options based on writer’s intention, the system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
Figure 3. Vistoria balances precision and exploration through two complementary generation modes: in Exact Craft mode, a single card closely follows the author’s intention to concretize ideas; in Creative Spark mode, three cards are generated with deliberate variations in characters, settings, or objects, offering diverse options to inspire new directions.
Figure 4. A set of visual instruments for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits. (b) Writers extract elements and compose across cards to discover new narrative directions, using collage alongside other instruments to recombine sketches, images, and text on the canvas. (c) Perspective Shift changes an image’s viewpoint and automatically regenerates the story’s point of view (first/third/second person). (d) Filters align visual style and textual tone (e.g., melancholic/dreamy) by jointly altering image effects and rewriting prose.
4.1.2. Synchronized Image–Text Co-Editing through Visual Instruments
Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination, yet existing tools separate these modalities and disrupt creative flow (DG2, DG3). To address this, we introduce a set of visual instruments (Figure 4) designed around three principles: (1) Reification, which draws on familiar operations from image editing (e.g., lassoing, collaging, filtering) and transforms them into persistent, manipulable operations that apply equally to text and image elements (Swerzenski, 2021; Masson et al., 2024a); (2) Polymorphism, which ensures the same operations apply uniformly across text and images; and (3) Structural Mapping, which synchronizes changes so that visual and textual representations carry the same narrative meaning. Together, these principles reduce modality switching costs while maintaining coherence between story segments and imagery.
Lasso. The lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text generates a new card focusing on the selected part with enriched narrative and visual details. Through polymorphism, the same selection logic applies across modalities—whether circling a visual detail or isolating a text segment—providing a consistent interaction pattern. The text within the selected area in the original content will be emphasized to form a new card. The extracted portion of text is used to regenerate the corresponding image. Structural mapping ensures that the lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 4 (a) ).
Collage. The collage instrument reifies the abstract act of “recombining inspirations” into a tangible manipulation: fragments of images, sketches, or text can be directly composed within a collage frame to form a new card. Through polymorphism, the same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation. Guided by structural mapping, the system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned. For instance, merging two character fragments not only produces a combined image but also generates a new story segment situating them together, ensuring that narrative and imagery evolve in sync (Figure 4 (b)).
Filter. Stylistic coherence is critical in fictional story writing, as consistent affective and aesthetic cues sustain narrative transportation (Green and Brock, 2000), activate readers’ interpretive schemas (Arbib, 1992), and enhance the emotional resonance of literariness (Miall and Kuiken, 1999). In Vistoria, the filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts both the style of the image and the tone of the accompanying prose (Appendix Table 4). Through polymorphism and structural mapping, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to act on both simultaneously. By making intangible stylistic intentions manipulable and synchronized, filters expand expressive possibilities while maintaining narrative immersion (Figure 4 (c)).
Perspective Shift. Fictional story writing often utilizes perspective shifts, as narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980). Cognitive poetics further shows that such shifts alter readers’ empathy and immersion. First-person narrations foster intimacy, while third-person perspectives enable broader structural awareness (Keen, 2007). The perspective-shift instrument reifies this narratological concept into an actionable operation: changing the visual viewpoint of a scene automatically regenerates the story fragment from a first-, third-, or second-person perspective. Through polymorphism, this instrument applies consistently across modalities, altering either an image or its accompanying text triggers a corresponding adjustment in the other. Structural mapping ensures that the focalization shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and structural awareness in a synchronized manner (Figure 4 (d)).
4.1.3. Highlight Elements and Cluster.
Writers often struggle to integrate scattered highlights and annotations on the Canvas into coherent storylines, leaving ideas fragmented across cards (DG4). Vistoria addresses this by transforming the dispersed fragments into reusable narrative building blocks, aligning them structurally across characters, objects, and scenes. This process is centered in the cluster panel (Figure 5), which turns fragmented inputs into organized knowledge assets.
Figure 5. Writers highlight objects and text segments on cards(b-1); the Cluster panel aggregates these by character/object/scene and can auto-summarize settings/plot/description about a certain object to guide final writing(b-2); Clicking on a specific image reveals the corresponding highlights and comments from earlier phases left on canvas(b-3).
On the canvas, writers can highlight textual segments, edit stories directly, and add inline comments noting potential uses in later drafting. Story objects, such as characters, settings, or scenes, are represented as editable keywords that can be highlighted by themselves. The system automatically binds each highlighted object to its associated text, consolidating references across multiple cards.
The cluster panel then aggregates all highlighted objects into an organized overview of evolving narrative elements. This eliminates the need to manually scan scattered cards and provides writers with a dynamically updated, object-centered workspace. Selecting an object reveals its complete associated materials, such as linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing.
Beyond simple aggregation, the panel supports higher-level knowledge construction through its summary feature. When this feature is invoked, the system generates structured summaries of settings, descriptions, and plot elements derived from highlights and comments. These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2022), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
4.2. Implementation
System Architecture. We adopted a decoupled front-end/back-end architecture. The React front-end enables efficient rendering for complex interactive interfaces, while the Flask back-end flexibly handles model calls with minimal overhead. Axios manages asynchronous communication between layers. The canvas module uses React-Flow for dynamic node-edge relationships and integrates “path to SVG” for natural interactions like freehand drawing and lasso selection. A DOM-based screenshot function ensures visual consistency in exported images. The text editor integrates React-Quill for rich text editing with lightweight customizations. Global state (text content, canvas nodes, selections) is centrally managed via Zustand to ensure consistency across modules. Per-session persistence relies on sessionStorage, which preserves data during the study session and clears on tab close to protect privacy.
Back-end Multi-Agent flow. The back-end bridges front-end actions with language and vision models through multi-agent orchestration. Every canvas action (e.g., selection, lasso, collage, inline text) is rasterized into screenshots paired with local context (outline, prior cards, global theme). A text agent (o4-mini) examines screenshots using templated prompts from a prompt library to extract user intent, recover sketch layouts, and write text consistent with surrounding outlines. The agent returns structured JSON (story text, intent, layout hints, semantic tags) that downstream visual agents use for synthesis and alignment.
Image Generation and Editing. An image agent invokes Flux Kontext Pro with consolidated prompts that fuse screenshot-derived intent, story fragments, and style controls. Sketches serve as structural scaffolds respecting composition while avoiding duplication; reference images act as content anchors, encouraging variation over copying. When neither is present, Flux operates in free mode guided by text-agent intent and style.For editing, front-end instruments (e.g., perspective shifts, filters, collage) are translated into prompt deltas through GPT-4o’s image pathway, yielding coupled visual updates. In parallel, the text agent revises tone or viewpoint for synchronized multimodal transformation. All artifacts are persisted as cards (image, story, objects keywords), enabling highlight-and-cluster operations that help reassemble fragments into coherent narratives.
5. User Study
To evaluate the usability, effectivness and usefulness of Vistoria, we conduct a lab user study with 12 participants. The study aims to answer the following research questions:
  • RQ1: What interaction patterns and strategies emerge when participants use a image-text co-editing process?
  • RQ2: How does image-text co-editing facilitate and shape the cognitive process of story creation, in terms of ideation, iteration, and refinement?
5.1. Participants
We recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience. Most held Bachelor’s or Master’s degrees, with backgrounds spanning science, arts, design, or communication. Their creative practices included fiction writing, screenwriting, songwriting, advertising, research, and philosophy. Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=2) or 4–7 years (n=3).
All participants were familiar with large language models (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, worldbuilding, and style imitation. This diversity reflects our design goal of supporting not only expert writers but also a broader range of users. Each participant received \$40 USD compensation.
5.2. Procedure
5.2.1. Apparatus
Sessions were conducted on a laptop computer with keyboard and mouse for typing, dragging, and selecting. To support sketch input, we provided an external tablet (iPad) for freehand drawing on the canvas.
The baseline condition presented a side-by-side interface with a text editor and GPT-4o conversational panel, enabling both manual editing and LLM-assisted text/image generation. Participants completed two story-writing tasks (Appendix A.1), each extending a given story beginning into a 300–500 word draft. Tasks were counterbalanced across conditions (Baseline vs. Vistoria).
5.2.2. Study Procedure
The study followed a within-subjects design with counterbalanced condition order. After informed consent and a demographic survey, participants were introduced to Vistoria through a written guide and tutorial video, followed by a short hands-on exploration (15 minutes).
In each condition, participants first focused on worldbuilding and idea exploration (20 min) and then on refining and improving the story (20 min). We divided the writing task into two phases (exploration and refinement) to prevent participants from prematurely committing to a single storyline and to reduce fixation, thereby encouraging broader ideation before focused improvement (Suh et al., 2024).
After each condition, participants completed surveys including NASA-TLX (Hart and Staveland, 1988), Creativity Support Index (CSI)(Cherry and Latulipe, 2014), self-perceived ML experience(Choi et al., 2024; Wu et al., 2022), and (for Vistoria only) perceived feature usefulness. All instruments used 7-point Likert scales. Following both conditions, participants completed a 15–20 minute semi-structured interview.
All sessions were video-recorded via Zoom. We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
5.2.3. Data analysis
We employed a mixed-methods approach to systematically analyze three types of data.
First, interview transcripts were independently open-coded by two authors following an inductive, grounded theory–informed process (Thornberg, 2012); initial codes were iteratively refined into a shared codebook, with themes developed through constant comparison. Disagreements were resolved via negotiated agreement, and analytic memos were maintained for transparency.
Second, interaction data were analyzed through structured video coding by two authors and aligned with system logs on an event-by-event basis. We examined the sequences of function use and compared exploration patterns across the baseline and our condition. To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance. When assumptions were violated, we used the Wilcoxon signed-rank test.
6. Study Results
Figure 6. Interaction records of all participants. The creative workflow begins with multimodal generation—primarily text, complemented by sketches or images to express intentions, followed by refinement and iteration using visual instruments, during which textual descriptions are continuously revised in parallel. Definition of specific behaviors: Generation includes multimodal creation of new cards using Creative Spark, Exact Craft, or Collage; image-based editing refers to operations such as Lasso, perspective shift, and filters; text-based editing covers modifications of generated story segments on the canvas as well as edits made in the text editor; other operations include updates through highlight, cluster, and upgrading global settings.
Vistoria Baseline Statistics
mean std mean std p Sig.
Transparent of ML model Transparency 4.250 1.865 3.417 2.234 0.2015
Controllability 4.750 1.357 4.833 1.586 0.8664
Sense of Collaboration 5.583 1.505 4.333 1.670 0.0206 *
Support for Thought Process 5.333 1.371 4.750 1.960 0.3172
Match to goal 5.500 0.905 4.833 1.337 0.1201
NASA-TLX Mental 5.16 1.528 3.167 1.337 0.0000 **
Physical 4.667 1.723 2.083 0.669 0.0002 **
Temporal 2.917 1.443 2.750 1.215 0.7723
Effort 4.083 1.443 3.500 1.834 0.3388
Performance 5.250 1.712 5.083 1.564 0.7986
Frustration 2.750 1.183 1.750 0.622 0.0204 *
Creativity Support Index Exploration 4.917 1.240 4.750 1.485 0.7126
Expressiveness 6.083 0.996 4.333 1.775 0.0232 *
Immersion 4.917 1.505 2.750 1.545 0.0006 **
Enjoyment 5.333 1.435 4.917 1.379 0.1753
Results Worth Effort 5.250 1.357 5.583 0.793 0.5166
Collaboration 5.500 0.674 4.583 1.505 0.0418 *
Table 2. Comparison of survey results: Vistoria vs. Baseline. Sig.: * $p<.05$; ** $p<.01$
6.1. Strategies for Integrating Multimodality in Creative Writing
To address RQ1, we identified four key interaction strategies employed when using Vistoria.
First, participants leveraged multimodal elements to externalize and interconnect different branches of their ideas. Second, they utilized image-text pairs as catalysts to generate richer and more dynamic narrative descriptions. Third, participants employed visual instruments as a mechanism to drive pivotal shifts in their storytelling. Finally, they treated images as both checkpoints and traces, enabling them to maintain narrative coherence and facilitate recall throughout the story development process.
6.1.1. Use Multimodality to Support Expressing and Connecting Ideas
Log data (Figure 6) and observations (Figure 13) reveal that participants followed a consistent creative workflow when using the system. The process typically began with multimodal content generation, where participants combined generation functions across multiple modalities to visualize their intended content. During this exploratory phase, participants iteratively collected and refined story elements through generated cards(images-text pairs), at the same time, frequently modifying text in the editor or regenerating images based on updated text.
As participants explored multiple visualization directions, they employed the Collage function to combine existing canvas content and discover new narrative possibilities, or used visual instruments (Lasso, Filter, Perspective Shifts) to simultaneously modify text and images. This process involved continuous comparison with LLM outputs to determine desired versus undesired elements (P1, P2, P4, P8, P10). Participants systematically highlighted useful images and text for final translation into the text editor, often using Cluster functions to summarize character or object settings for incorporation into their final writing.
Multimodal Expression Enhances Creative Intention. The multimodal integration enhanced participants’ ability to convey creative intentions. As shown in Figure 6 (b), text served as the primary medium but was consistently supplemented with sketches and images for context, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs. $M_{\text{Baseline}}=4.33$, p = .023) (Table 2). P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.” (Figure 11). Nearly all participants (P2, P6, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions. P7 illustrated this with a concrete case: she sketched a rough flower and archway, then added text specifying a glowing flower and a Gothic gate. The system fused the spatial layout from the sketch with textual details to generate multiple fitting images. She highlighted the unique value of sketching: “Drawing is pretty cool. In GPT, I want to draw a mental image, but GPT cannot… the geometry is always different from what’s in my head.“ (Figure 7). This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise creative outputs.
Figure 7. Multimodal Expression. Example where sketch structure + textual details jointly yield multiple relevant images; participants valued sketches to visualize spatial layout beyond what text-only tools could provide.
Collage Function Enables Creative Recombination. The Collage function emerged as a central strategy for creative recombination, receiving positive usability ratings (M=5.33, SD=1.5) (Figure 8). All participants used collage to merge extracted objects or scenes from different images, building connections across disparate elements. For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs. He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story (Figure 9). P11 articulated the creative freedom this technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” This multimodal recomposition enabled participants to quickly express envisioned scenes (P2) and provided “more freedom to envision and create the story” (P10). These practices highlight that collage is not merely a usability feature but a catalyst for multimodal recomposition, enabling participants to externalize, reconfigure, and expand their mental imagery into coherent narrative possibilities that text or images alone could not achieve.
Furthermore, multimodal input also increased participants’ immersion with Vistoria compared to the baseline ($M_{\text{Vistoria}}=4.92$ vs. $M_{\text{Baseline}}=2.75$, p = .0006) (Table 2). This heightened immersion stemmed from the playful, enjoyable nature of multimodal creation and the emotional reinforcement of freely expressing intentions across modalities. As P1 reflected: “Using the tool feels more like… doing something on a whiteboard or a big piece of paper, where you can do almost anything—it’s very free and interesting.” Together, multimodal inputs and outputs fostered a more engaging and emotionally resonant creative experience.
Figure 8. The perceived usefulness of each key functionality in Vistoria
6.1.2. Instrumental Interaction Tools Drive Narrative Development
Analysis of interaction logs (Figure 6 and 13) shows that participants often engaged with visual instruments in the mid-phase of creation, and these instruments directly shaped how stories were developed and revised.
Perspective Shift as a narrative frame-shifter.
The Perspective Shift instrument, which altered both the image and the story perspective, was rated among the most useful features (M=5.92, SD=1.56) (Figure 8). P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework. Changing the perspective of the voice of the story added personal sentiment and changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 10 (a)). Perspective Shift allowed authors to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
Lasso as a granularity controller for local-to-global rewriting.
The Lasso instrument was also valued as useful (M=5.0, SD=1.28) (Figure 8) for enabling writers to zoom between different narrative scales. P8 emphasized, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.”(Figure 10 (b)) This reflects how the lasso tool enables a narrative “zoom” functionality, allowing writers to oscillate between macro-level story development and micro-level detail refinement within a single interface. Similarly, P9 and P11 described how the lasso allowed them to deepen or refine specific text and emphasize the key points with more focused attention. In this way, the Lasso operates as a narrative instrument, turning macro-level edits into micro-level adjustments while preserving precision, enabling rollback, and sustaining fluent exploration.
Figure 9. Collage. Participants used Collage for creative recombination, merging extracted objects/scenes (often mixing sketches with images) to specify visualization and advance story ideas; usability was positively rated.
Filters as affective parameterization for tone alignment.
The Filter instrument shaped narrative emotion and tone by visually parameterizing affect. P11 noted, “The image provides the style, which influences my story’s tone and direction… After using the filter, I can’t determine tone from text alone—I need to choose between suspenseful or romantic, and the visuals help me decide.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with filters, and the girl suddenly turned into a sad face. It’s very easy to see what it’s doing.” (Figure 10 (c)). The immediate visual feedback fostered a more intuitive workflow, aligning emotional intent with visual representation. The immediate visual feedback aligned emotional intent with representation, streamlining tone-setting decisions.
Taken together, these tools acted as narrative instruments that transformed localized operations into meaningful shifts in viewpoint, scale, and tone. They enhanced authors’ control, amplified expressiveness, and sustained a fluid perception–action loop throughout the creative process.
Figure 10. (a) Perspective Shift was rated highly useful (M=5.92); changing the image viewpoint also reframed narrative voice and redirected story development, as participants reported; (b) Filters (M=4.25) synchronize mood and style across media—applying a visual filter also rewrites the associated text to match the intended emotional tone, ensuring stylistic coherence; (c) Lasso tool (M=5.0) enabled writers to focus at different narrative scales by extracting or isolating elements to steer local and global edits.
6.1.3. Structural Mapping between Images-Text Paris as a Catalyst for Vivid Story Writing
Participants emphasized that when text and images appeared together, images served as direct references during writing, easing detailed description and enriching narrative depictions. This integration worked through two cognitive mechanisms that supported vivid story construction.
Images-Text Paris as cognitive scaffolding for unfamiliar scenarios. Visual references helped participants imagine actions or settings beyond their lived experience. As P11 reflected, “I’m more comfortable describing conversations and overall scenarios than people’s actions—especially kids, who might have more interesting reactions. Visualizing with text descriptions helps me imagine unfamiliar actions, which I can then use in my story.” Here, the image acted as a structural anchor by aligning the visual cues of children’s postures and expressions with possible textual descriptions. In other words, structural mapping between what was seen (visual structure) and what could be narrated (linguistic structure) scaffolded richer and more plausible action descriptions.
Images-Text Paris enabling descriptive language translation. Participants also converted visual elements into vivid prose, effectively mapping structures across modalities. As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the final images in her writing and directly adopted text from the generated cards, integrating them into her final story. (Figure 12) This visual-to-textual translation was also evident in concrete writing outcomes. For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text such as “Claire steadies the box on her lap” and “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story. (Figure 11) Here, the structural alignment between gesture in the image and embodied description in text allowed P1 to carry emotional undertones across modalities, turning static imagery into dynamic, sensory-rich narrative moments.
Overall, structural mapping between visual and textual representations scaffolded descriptive richness, enhanced vividness, and supported schema formation for scenarios participants might otherwise have struggled to imagine.
Figure 11. Visual-to-text translation: writers turn visual cues into vivid prose that readers can picture; sample lines show how P1 used sensory-rich descriptions derived from an image-text pair.
6.1.4. Structural Mapping across Multiple Dimensions to Trace the Story
Participants used not only text but also images as visual checkpoints to trace story direction and development. Because images were more immediately perceptible than text, they helped participants align outputs with intended style or mood and, when viewed in sequence, detect progressions and gaps (P6, P7, P2). Within Vistoria, this process evolved into structural mapping: participants systematically compared visual–textual structures across dimensions, gaining multimodal visibility into narrative development.
This layered view created a relational space where main storylines and branching alternatives could be mapped against one another, allowing participants to perceive both continuity and divergence, thereby fostering coherence while supporting flexible redirection. As P12 noted, visuals alongside text set the mood and made it easier to recall earlier ideas, illustrating how structural mapping provided mnemonic scaffolding that supported continuity in storytelling. Compared to GPT’s linear text-only workflow, participants found that Vistoria’s visual traces better supported narrative management and reduced idea drift or loss (P12, P11, P2). P2 reflected on this benefit explicitly:
“Using the system, we started with a baseline picture and then came up with another picture… it was easy to trace the development of the story. If you get too far down that chain and don’t like it, you can just delete that node and go in a different direction. I liked visually being able to see the progress from generation to generation.”
Here, structural mapping across successive images enabled participants to align branches and selectively prune or redirect them. Such multi-dimensional mapping clarified narrative progression, improved recall and coherence, and supported flexible branching and gap detection, sustaining sensemaking throughout the evolving story.
Figure 12. Images act as cognitive scaffolds, helping writers describe unfamiliar actions or contexts more concretely. P5 constructed her final writing by referring to both images and adopted and edit related text.
6.2. Multimodal Approaches Creation to Inspire Creativity
To answer RQ2, we found that image-text co-editing facilitated creativity in fictional story writing through three key cognitive processes: (1) divergent exploration, where participants spatially arranged multimodal components and pursued multiple narrative directions; (2) serendipitous discovery, where unexpected LLM outputs sparked new creative possibilities; and (3) multimodal cues activating memory and cultural connections. Together, these processes enhanced participants’ sense of authorship by positioning the system as a collaborative partner that expanded creative possibilities rather than replacing human imagination.
Figure 13. Behavioral diagram contrasting Vistoria vs. baseline: participants cycle through multimodal generation, collage/recombination, and coupled image–text editing before collecting highlights for integration. Data from P8-P12 show that, compared with baseline, participants using Vistoria explored more directions, with greater divergence (Branches) within each direction. Participants also tended to pursue deeper exploration within specific directions when using Vistoria.
Vistoria Baseline
Mean # of directions 6.92 $\pm$ 2.81 1.42 $\pm$ 1.08
Mean # of branches 3.00 $\pm$ 1.35 1.92 $\pm$ 0.67
Mean depth 1.70 $\pm$ 1.18 2.00 $\pm$ 1.22
Table 3. Descriptive statistics (mean $\pm$ SD) for Vistoria vs. Baseline. When using Vistoria, participants exhibited broader exploration; at the same time, as shown in Figure 13, they also tended to pursue individual directions with greater depth. Specifically, Directions denote the number of distinct aspects or dimensions explored when co-creation with Vistoria or baseline. Branches represent the diversity of possibilities generated within a given direction. Depth indicates the mean number of iterative steps within each branch.
6.2.1. Bottom-up Creation to Support Divergent Exploration
As shown in Figure 13 and Table 3, participants demonstrated significantly greater breadth and depth of exploration with Vistoria compared to the GPT baseline. They pursued a wider range of narrative directions ($M_{\text{System}}=6.92,SD=2.81$ vs. $M_{\text{Baseline}}=1.42,SD=1.08$) while also producing more variations within single paths (branches: $M_{\text{System}}=3.00,SD=1.35$ vs. $M_{\text{Baseline}}=1.92,SD=0.67$). Although the average depth per path was slightly lower for Vistoria ($M_{\text{System}}=1.70,SD=1.18$ vs. $M_{\text{Baseline}}=2.00,SD=1.22$), this reflected participants’ substantially broader exploration overall; within individual directions, they often engaged in deeper iterations, as evident in Figure 13. Such enhanced exploration transformed their creative process from passively consuming LLM outputs to actively curating multimodal components.
The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints. This freeform environment encouraged spatial arrangement, clustering, and repositioning of ideas through manipulation and recombination of alternative story–image cards to explore divergent scenarios (P7, P4, P5, P8). Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases, positioning the system as an expressive medium rather than merely a text generator. This exploratory approach helped participants avoid early fixation and sustained creative engagement. As P8 observed: “GPT workflow is more streamlined… top-down. Using the visual app feels more bottom-up. You are open to possibilities, so there’s not a finite result and more possibilities being explored.” Ultimately, this shift highlights how multimodal, spatially organized exploration fosters divergent thinking, sustains creative momentum, and supports a more generative form of narrative authorship.
6.2.2. Serendipity from LLM Randomness Enriches Narrative Development
Another recurring theme was that the inherent randomness of LLM-generated image-text output often introduced serendipitous elements into participants’ creative process. Rather than being perceived as noise, these unexpected details were frequently embraced as inspiration, enriching story development and occasionally opening entirely new narrative directions.
For example, when P1 generated an image-text pair of the mysterious box, the system unexpectedly inserted the logo “LEH.” on the box and in the text description. Intrigued, P1 subsequently asked the system to visualize who or what “LEH” could be, gradually evolving this accidental motif into a recurring symbol within their final story (Figure 11). Nearly all participants described how surprises in LLM-generated images or texts gave them fresh possibilities to explore. P12 captured this dynamic vividly: “The system gave me a riddle at an early stage, I didn’t understand it at first, but as I kept working, I slowly decoded it, and this process of unraveling became part of my story. It helped me construct something unique that I wouldn’t have created otherwise.” (Figure 14) Together, these examples show how the interplay of multimodal generation and LLM’s inherent unpredictability fueled creativity. Randomness supported divergence by surfacing novel ideas, while participants retained agency in selectively integrating these elements, leading to unique, personally authored narratives.
Figure 14. In the creation process of P12, unexpected LLM randomness served as serendipitous prompts: a generated riddle was decoded and transformed into a key-and-note motif guiding Claire toward the lighthouse, illustrating how stochastic details became generative cues that enriched development and redirected the plot.
6.2.3. Unlocking Creativity Through Memory and Cultural Resonance
Participants reported that multimodal cues routinely triggered personal memories and cultural references, enriching their stories. Visual props, e.g., a medieval mask, readily invoked film/literary associations; as P4 noted, “Seeing the mask immediately brought Batman/Phantom to mind… it made choosing a style much easier.” (Figure 15) Crucially, image–text pairs co-triggered lived experiences. In P7’s case, a fountain scene image together with the phrase “mossy stone” jointly evoked a vivid memory, which she then translated into a concrete narrative element, the envisioned fall that became her story’s ending(Figure 15). These multimodal traces turned associative recall into usable story material, weaving authenticity and imaginative depth into the narrative.
6.2.4. Strengthened Authorship and Agency
Enhanced exploration with Vistoria strengthened participants’ sense of authorship. While GPT outputs in the baseline condition were consistently described as “surface-level” (P2, P5, P7) and felt like “someone else’s work” (P3), almost all participants reported markedly stronger authorship with Vistoria, enabling deeper exploration and richer plots (P1, P2, P3, P5, P12, P10). P5 articulated this shift:
“When using Vistoria, every idea originated from my own imagination, and the final story was formed by combining these different self-generated ideas. This gave me a strong sense that the story was truly my own creation. I think of the system less as a generator and more as a companion sketchbook. While you explore different directions, you always remain in charge of the final piece.”
P9 reinforced this sentiment: “This tool is more ‘me’… I control characters and plots; ChatGPT may not become my story.” Quantitative results corroborated these perceptions: participants rated Vistoria as providing stronger collaborative experiences (CSI: $M_{\text{Vistoria}}=5.50$ vs. $M_{\text{Baseline}}=4.58$, p = .0418) (Table 2). Participants framed Vistoria as a supportive co-pilot rather than a replacement; as P10 noted: “I consider this system as a copilot that does not help you with everything, and this system that you have to create by yourself.” Together, these findings mark a transition from passively adopting others’ ideas to actively curating one’s own generative outputs, reinforcing creative agency and ownership.
Figure 15. Evolving cards leave visual traces that cue personal memories and cultural references (e.g., a mask → Batman/Phantom; a “mossy stone” → place memory), which writers externalize into concrete scenes. These multimodal traces act as mnemonic scaffolds, helping retrieve earlier thinking and weave lived experience into the story world—transforming associative recall into narrative material.
6.3. Tradeoffs of Multimodal Interaction for Creative Writing
While image–text co-editing enhanced authorship, coherence, and expressiveness, it also increased workload. NASA–TLX scores showed significantly higher mental demand ($M_{\text{Vistoria{}}}=5.16$ vs. $M_{\text{baseline}}=3.17$, p=.0000) and physical demand ($M_{\text{Vistoria{}}}=4.67$ vs. $M_{\text{baseline}}=2.08$, p=.0002), plus moderately higher frustration ($M_{\text{Vistoria{}}}=2.75$ vs. $M_{\text{baseline}}=1.75$, p=.0204) (Table 2). These increases reflect the effort of coordinating across modalities and actively curating outputs relative to the GPT baseline. Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10): “The biggest burden is switching between tools to sketch or type; it takes time to learn and adapt, even though the functions are useful.” Additional difficulty arose from unfamiliarity with the canvas interface (P3, P10).
6.3.1. Higher burden of authorship.
Almost all participants valued the ability to maintain control of the story, and the creation process gave them a stronger sense of ownership. Yet, this empowerment came with a cost: several participants noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9). Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system required users to supply and elaborate on their own ideas before meaningful generation occurred, which may lead to a higher mental workload.
This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply receiving GPT’s ready-made text (P1, P2, P11). Thus, stronger authorship agency came at the cost of higher workload.
6.3.2. Validating ideas rather than generating them.
Some participants noted that the system’s strengths lay in validating or expanding concrete mental images rather than generating abstract directions(P1, P11). As P1 explained, “when I have a vague impression in my mind, I tend to generate some image-text pairs. But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities. In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration.
Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and offering prompt-like guidance, whereas the multimodal Vistoria system primarily served to elaborate or diverge from existing visions. This made the multimodal approach particularly valuable for writers with partially formed concepts, but less helpful during the most open-ended phases of ideation when abstract exploration is more important than visual specificity.
6.3.3. Cognitive offloading and efficiency despite higher load.
Despite a higher mental workload, participants repeatedly emphasized that multimodality enabled more efficient allocation of attention. From participants’ view, the image-text pairs externalized fleeting ideas, preserved spatial and sensory detail, and reduced information loss when translating imagination into tangible artifacts (P4, P7). As P7 noted, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.”
Thus, while Vistoria required more effort and decision-making, it also freed participants from lower-level memory maintenance and allowed them to focus on higher-order creative synthesis. With greater familiarity, the efficiency gains from this offloading may outweigh the initial overhead.
7. Discussion
7.1. Cross-modal LLM-Mediated Instrumental Interaction
We extend the notion of Instrumental Interaction into an LLM-mediated, cross-modal paradigm where generated artifacts repeatedly re-enter the workflow as manipulable outputs. In classic formulations, instruments mediate between users and domain objects, maintaining a separation between “tool” and “content” (Beaudouin-Lafon, 2000). In generative contexts, this dichotomy breaks down: each card (image–text pair) is simultaneously an outcome and a renewed object of operation, enabling recursive, non-linear composition (Riche et al., 2025).
This reframing also reshapes reification. Vague creative intents expressed through text, sketches, or images materialize across modalities into manipulable, cross-linked artifacts that remain editable and reusable. Reification thus spans modalities rather than existing in a single channel, supporting direct manipulation and reflection on both structure and semantics of evolving narratives (Riche et al., 2025; Masson et al., 2024b).
A key mechanism enabling this circulation is structural mapping (Gentner, 1983; Norman, 2013; Masson et al., 2024b). Participants made sense of their evolving materials not only by inspecting the surface properties of each medium but also by exploiting the correspondences in shared properties, concepts, or identities established between them. Refining a visual scene, for example, carried implications for narrative structure, tone, or pacing in text, while textual edits translated into compositional adjustments in images. In this way, mappings transformed what is usually a cognitive act of alignment into an operational resource: they stabilized coherence while keeping alternatives visible and accessible. Instrumental interaction and structural mapping thus operate as mutually reinforcing processes. Instruments generate the artifacts that mappings align, while mappings scaffold the interpretability and reuse of instruments.
This dynamic carries several design implications for future systems. They should not only offer isolated multimodal tools but cultivate instrumental ecosystems in which mappings are first-class and inspectable, and where families of instruments carry coherent semantics across domains (Riche et al., 2025). Furthermore, instead of overwhelming users with immediate complexity, operations can be gradually disclosed, allowing mappings to surface progressively and adapt to context (Shneiderman and Maes, 1997; Allen et al., 1999). Provenance views that expose the cascade of edits across modalities could support both divergence and recombination, helping individuals and teams understand how narratives branch and converge (Masson et al., 2024b).
Ultimately, the LLM can be positioned less as a generator of content and more as a mediator of instruments that align structures, propagate semantics, and synthesize new operations from prior materials. This shift reimagines authorship as an orchestration of evolving, cross-modal instruments, opening a design space where exploration and coherence can be sustained simultaneously (Sawyer, 2017; Masson et al., 2024b).
7.2. Orchestration and Curatorial Authorship to Form Creativity in Cross-Modal Creation
Traditional creativity theories emphasize individual expression through language and pre-formed intent (Smith, 2020; Donnelly, 2011), positioning the creator as the primary originator of content. Our findings challenge this view by demonstrating that creativity in Vistoria manifested as orchestration. Participants conducted and harmonized streams of multimodal LLM outputs through selection, combination, and reframing. Creative goals emerged through interaction with generated materials rather than preceding them, with meaning composed across representational boundaries. This suggests creativity in AI-augmented contexts should be theorized as the strategic coordination of evolving possibilities, repositioning authorship from origination to curation and synthesis within a collaborative ecosystem.
While existing multimodal creativity research often focuses on seamless integration between modes (Yen et al., 2025; Rosenberg et al., 2024; Shi et al., 2025), our study reveals that productive tensions between text and image became central to the creative process. Participants actively negotiated the trade-offs between imagistic specificity and narrative openness, treating these tensions not as problems to solve but as resources for exploration. The ongoing adjudication process—accept, adapt, defer, or discard–became the primary engine of creative coherence. This extends theories of creative constraint by showing how representational tensions across modalities can drive rather than hinder creative development.
Furthermore, our findings contribute to debates about human agency in LLM-assisted creation by demonstrating that curatorial work enhanced rather than diminished participants’ sense of ownership. Authorship was experienced as steering a dialogue with the model, including deciding what counts, where elements belong, and how disparate fragments accrue into a unified voice. This sense of agency grew precisely from having meaningful choices to negotiate among viable alternatives (Resnick et al., 2005). This suggests a new model of distributed authorship where human control emerges through deliberate curation within a space of LLM-generated possibilities.
As for design implications, our findings suggest that future system design should prioritize support for orchestration over seamless, one-shot generation. Instead of producing singular outputs, tools could by default surface multiple cross-modal alternatives, accompanied by lightweight mechanisms that enable users to accept, adapt, defer, or discard options as part of an ongoing multimodal curatorial process (Resnick et al., 2005). Moreover, systems should not only present outputs but also make the inherent tensions between modalities visible and adjustable. This can be done, for example, through interactive controls such as ambiguity sliders that allow users to navigate between imagistic specificity and narrative openness.
Beyond simple interface controls, future systems could also provide deeper support for curatorial authorship. For example, they might keep track of where each fragment or idea came from across different iterations, or offer reusable “curation moves” that let users quickly repeat and adapt common ways of combining or editing materials (Shneiderman, 2007). In this way, design emphasis shifts away from achieving frictionless integration and toward helping users actively steer evolving creative possibilities, foregrounding authorship in AI-mediated contexts as a matter of curation, synthesis, and negotiated choice (Ding, 2024).
7.3. Limitation and Future Work
While our study provides valuable insights into multimodal story writing, several limitations constrain the generalizability and scope of our findings.
Task Scope and Short-Term Focus. The 300–-500 word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures. We did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over the time of use. Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
In the future, we plan on conducting field studies to deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods. Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles. We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization. These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
Participant Sample. Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups. Furthermore, the participant pool has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented. Future research should address these limitations by recruiting a larger and more diverse set of participants, including professional authors, older and younger writers, and individuals from varied cultural backgrounds, to more fully evaluate the system’s applicability and generalizability.
8. Conclusion
This paper introduced Vistoria, a multimodal text–image co-editing system designed to support the cyclical “planning–translation” process in fictional writing. Grounded in Wizard-of-Oz co-design study and theories of Instrumental Interaction and Structural Mapping, the system integrates four key capabilities: (1) representing sketches, texts, and references as manipulable card units; (2) enabling cross-modal alignment so that edits in one modality propagate to the other; (3) providing polymorphic operations that apply consistently across text and image; and (4) supporting organization and reuse through highlighting and clustering panels. A controlled study with 12 creative writers demonstrated that Vistoria enhanced writers’ expressiveness, immersion, and sense of collaboration, while broadening exploration of narrative branches and reinforcing authorship and control. These findings suggest that Vistoria transforms the writing process into a dialogic interaction with multimodal instruments, enabling authors to negotiate meaning and control across modalities.
Appendix A Appendix
A.1. Writing Topics Used in the User Study
The two writing topic prompts used in the user study were:
Topic 1—Claire steps outside her apartment and finds a small wooden box on her doorstep. The box is secured with an old brass clasp and feels unexpectedly heavy when she lifts it. Its surface is scratched, as if it has been handled many times, and faint traces of dried sea salt cling to the edges.
Topic 2—During her morning jog through the park, Maya discovers an ornate iron gate hidden behind overgrown ivy. Through the bars, she can see a path lined with luminescent flowers that pulse gently like soft heartbeats. The air carries faint whispers in a language that sounds hauntingly familiar, almost like someone calling her name.
A.2. Filters
The following are the types of filter supported by the filter instrument, showing how different types of filters are applied to image styles and mapped to text tone or emotion.
Filter Image Effect Text Effect
Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasizes positivity, vitality, intimacy
Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflects contemplative and stable moods
Dramatic Deep blacks, sharp whites, directional lighting $\rightarrow$ create intensity, mystery, urgency Heightens stakes and emotional tension
Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Supports subtle, nostalgic, introspective narration
Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopts reflective and universal tone
Table 4. Filter types with corresponding image and text effects.
A.3. LLM Prompts
Communication between different agents is carried out through JSON-formatted messages, ensuring efficient and reliable information transfer.
All prompts are structured in the following format:
  • Blue text indicates placeholders that will be replaced with task-specific information in different generation requests.
  • Pink text represents the mandatory output type of information generated by the agent.
The below is an example prompt for precise description generation of GPT-4o.
You are a visual story developer who analyzes screenshots to decode user visualization intent and creates detailed story segments that bring their creative vision to life. Process: 1) Examine the screenshot to understand what specific story content the user wants generated by identifying: printed text({text} ) which is the primary indicator of what the user wants you to generate, handwritten text expressing the user’s desired content direction and story focus where generation should address gaps and missing details, any images or illustrations with reference text {previous_text} for additional context, and hand-drawn sketches representing scenes from the user’s imagination. Synthesize these elements to understand the user’s envisioned story. The generated story should mainly focus on filling in content not covered in ({text} ) instead of still remain unknown. 2) If hand-drawn illustrations exist in the screenshot, return the information about the story scene conveyed in the illustration’s layout; if none exist, output ’none’. 3) Generate Focused Story Content: Using the existing written passages {full_text} only as background context to ensure logical consistency, create a NEW detailed story segment that elaborates on a specific scene or moment the user wants to visualize, focuses primarily on the intent expressed in the screenshot rather than expanding the existing text, contains concrete details, character emotions, environmental descriptions, dialogue, and interactions, maintains consistency with the global theme {global_theme}, and can reference previous text {previous_text} if relevant to the visualization goal. 4) The narrative should contain substantial plot or setting content, not just descriptive language. The generated story should introducing new, insightful elements based on the context and provide new direction of the story development that can reificate the story. The generated stories need to be imaginative with concrete content, not filled with uncertainties. Respond in JSON format: { "story": "A detailed paragraph of no larger than 100 words that creates NEW story content focused on the user’s screenshot intent, elaborating a specific scene with concrete details while maintaining logical consistency with the background context.", "intention": "The visualization intention read from the screenshot for story generation direction", "sketch_information":"Regarding line sketches, integrate them with story descriptions to capture the user’s envisioned layout and scene details communicated through hand-drawn imagery, directing the image generator to produce story scenes based on this layout guidance. Avoid generating stories that may trigger content moderation." }. Provide only the JSON response without markdown formatting or additional commentary. Let’s think step by step.